9. A solid sphere of mass M, radius R and having moment of inertia about an axis passing through the centre of mass as I, is recast into a disc of thickness t, whose moment of inertia about an axis passing through its edge and perpendicular to its plane remains I. Then, radius of the disc will be (A) 2R/√15 (B) R√(2/15) (C) 4R/√15 (D) R/4

9.     A solid sphere of mass M, radius R and having moment of inertia about an axis passing through the centre of mass as I, is recast into a disc of thickness t, whose moment of inertia about an axis passing through its edge and perpendicular to its plane remains I. Then, radius of the disc will be


        (A) 2R/√15


        (B) R√(2/15)


        (C) 4R/√15


        (D) R/4

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mainak

moment of inertia of solid sphere I =2/5 MR2

mass will remain conserve

 let the new radius is R1

so I = MR12/2 + MR12

       =3/2 MR12

 

so 3/2 MR12 =2/5 MR2

    R1 =2R/√15

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin